OpenStudy (anonymous):

solve the simultaneous equation 5x + 6y =37 2x + 3y =4

OpenStudy (anonymous):

multiply 2x+3y+4 with 2

OpenStudy (anonymous):

5x+6y=37 4x+6y=8

OpenStudy (anonymous):

now subtract both of them u get

OpenStudy (anonymous):

x=29

OpenStudy (anonymous):

y=-18 x=29 2(5x+6y=37) ==10x+12y=74 ---------(A) 5(2x+3y=4)==10x+15y=20 ----------(B) A-B = -3y = 54 = = y = -18 put y= -18 in eq B x= 29

OpenStudy (anonymous):

5x + 6y = 37 2x + 3y = 4 -->(-2)2x + 3y = 4 ----------- 5x + 6y = 37 -4x - 6y = - 8 (result of multiplying by -2) ------------add x = 29 now sub 29 in for x in either of the original equations... 2x + 3y = 4 2(29) + 3y = 4 58 + 3y = 4 3y = 4 - 58 3y = -54 y = -18 check... 5x + 6y = 37 5(29) + 6(-18) = 37 145 - 108 = 37 37 = 37 (correct) ANSWER : x = 29, y = - 18 or (29,-18)